OpenStudy (anonymous):

Write an equation of a line in point slope form that has a slope of -3 and passes through the point (3, -4). Do not use spaces and enter fractions as improper reduced fractions when needed.

OpenStudy (lgbasallote):

this is pretty straightforward...plug it in the formula for point slope \[\huge y - y_1 = \frac{y_2 - y_1}{x_2- x_1} (x-x_1)\]

OpenStudy (lgbasallote):

oh wait...wrong formula

OpenStudy (lgbasallote):

use \[ y- y_1 = m(x-x_1)\]

OpenStudy (cwrw238):

use the formula y-y1 = m(x - x1) where m = slope and (x1,y1) is a point on the line

OpenStudy (anonymous):

its Y=mX+B

OpenStudy (lgbasallote):

nooo...that's the slope-intercept :)

OpenStudy (anonymous):

You're so quick lgba...I can't keep up lol

OpenStudy (anonymous):

ooh wait nevermind.. can someone show how to do it Step-by-Step please

OpenStudy (zepp):

I have a nice tutorial on that, can't find it :(

OpenStudy (lgbasallote):

lol you should see us back in the days =))))

OpenStudy (lgbasallote):

okay @Trexy what is the y-coordinate in this problem?

OpenStudy (anonymous):

@Calcmathlete wat do you think Hes smart.. and you too .. thats why I need help you guyz

OpenStudy (lgbasallote):

@zepp maybe you should make a table of contents too? :)

OpenStudy (anonymous):

Didn't eyad make a tutorial on this like just now?

OpenStudy (anonymous):

i didnt see it @Calcmathlete and please help me how to do this

OpenStudy (zepp):

AH!

OpenStudy (zepp):

>>>>>>>>>> Situation 3. The slope and a random point of the function are given I roll a dice.. I got 2 and 92, so our random point is at (2,92) I roll a dice again... got 3.. D: Point-slope form! \[\large y - y_{1} = m(x-x_{1})\]\[m= slope; ~~~~x_{1} ~and~y_{1}: A~point~on~the~function\] 2 is x1, 92 is y1 \[\large y - y_{1} = m(x-x_{1})\]\[\large y - 92 = 3(x-2)\]\[\large y - 92 = 3(x-2)\]Distribute...\[\large y - 92 = 3x-6\]\[\large y = 3x-6+92\]\[\large y = 3x+86\]

OpenStudy (lgbasallote):

like i said @Trexy what's the y-coordinate in the problem?

OpenStudy (zepp):

I kept the LaTeX version of my tutorial, this is an example :)

OpenStudy (anonymous):

lol @zepp maybe reread the problem?

OpenStudy (zepp):

Simply plug \(x_1,y_1\) and simplify it.

OpenStudy (zepp):

Calc, that was an example :x

OpenStudy (anonymous):

I know, but we don't need to put it into slope-intercept, just point-slope

OpenStudy (lgbasallote):

i think he was just quoting himself

OpenStudy (zepp):

:(

OpenStudy (anonymous):

Thank You .. I heart Everyone Here since I cant give all of you a Medal.. hahah

OpenStudy (anonymous):

ALright. @Trexy in order to use point-slope form, do the following. \[y - y_{1} = m(x - x_{1})\]Plug in the slope which is given and plug in the coordinates. m = -3 \(x_{1} = 3\) \(x_{2} = -4\)

OpenStudy (zepp):

\(y_2=-4\)*

OpenStudy (anonymous):

Oops Yeah. What ^ said.

OpenStudy (lgbasallote):

hmm i've been trying so hard since a while ago to teach but no ones responding lol..nevermind :p

OpenStudy (anonymous):

Me @lgbasallote ??